OpenStudy (anonymous):

Mr. Cruz is 40 yrs. old and his son is 12 yrs. old. How long will it take for Mr. Cruz to be as thrice as old as his son?

OpenStudy (ajprincess):

Let us say after x yrs Mr.Cruz's age becomes thrice the age of his son. then (40+x)=3(12+x) 40+x=3*12+3*x 40+x=36+3x Can u find x @punyeta?

OpenStudy (anonymous):

thanks.

OpenStudy (jkasdhk):

x=2 :)